# Navigating Public Sentiment in the Circular Economy through Topic Modelling and Hyperparameter Optimisation

### v1

@@ Abstract
To advance the circular economy (CE), it is crucial to gain insights into the evolution of public sentiments, cognitive pathways of the masses concerning circular products and digital technology, and recognise the primary concerns. To achieve this, we collected data related to the CE from diverse platforms including Twitter, Reddit, and The Guardian. This comprehensive data collection spanned across three distinct strata of the public: the general public, professionals, and official sources. Subsequently, we utilised three topic models on the collected data. Topic modelling represents a type of data-driven and machine learning approach for text mining, capable of automatically categorising a large number of documents into distinct semantic groups. Simultaneously, these groups are described by topics, and these topics can aid in understanding the semantic content of documents at a high level. However, the performance of topic modelling may vary depending on different hyperparameter values. Therefore, in this study, we proposed a framework for topic modelling with hyperparameter optimisation for CE and conducted a series of systematic experiments to ensure that topic models are set with appropriate hyperparameters and to gain insights into the correlations between the CE and public opinion based on well-established models. The results of this study indicate that concerns about sustainability and economic impact persist across all three datasets. Official sources demonstrate a higher level of engagement with the application and regulation of CE. To the best of our knowledge, this study is pioneering in investigating various levels of public opinions concerning CE through topic modelling with the exploration of hyperparameter optimisation.
@@ 1 Introduction
The concept of the Circular Economy (CE) has received increased attention from both academics and practitioners in recent years. The primary objectives of the CE are commonly regarded as achieving the twin goals of economic prosperity and improved resources efficiency [1]. Despite this, the impacts of CE on social equity and public perceptions are often overlooked or inadequately addressed. The implementation of CE also plays a critical role in the transition towards net-zero emissions [2, 3, 4, 5, 6]. By reducing resource consumption and optimising waste management, CE contributes to lowering energy demand and decreasing greenhouse gas emissions [7]. This relationship is especially significant in fields such as energy and artificial intelligence, where the integration of advanced technologies and circular principles can lead to substantial improvements in resource efficiency and sustainability.
@@
Moreover, public sentiment and understanding of CE remain notably limited. CE represents a paradigm shift from the traditional linear economy—characterised by the “take–make–waste” model [8] to one that promotes the principles of regeneration, material reuse, waste reduction, and pollution prevention. In a world grappling with pressing environmental challenges, public discourse on CE plays a pivotal role in shaping policies, business strategies, and societal perceptions of sustainable development. Therefore, it has become increasingly crucial to comprehend and interpret public sentiments concerning the CE. This urgency stems not only from the necessity to understand the intricacies of evolving public attitudes but also in the imperative to in-depth analyse the cognitive pathways that influence the public’s perception of recycled products within the broader framework of the CE [9].
The novel contributions of this paper are summarised as follows. Firstly, we conducted comprehensive and structured research based on our established framework. Our data collection efforts encompassed a wide range of sources, including posts from widely-used social media platforms such as Twitter and Reddit, where individuals express their views and thoughts on various topics, including the CE. Secondly, we analysed content from prominent news publishers like The Guardian, recognising their crucial role in disseminating information as well as shaping public discussions and influencing the topics of conversation. Thirdly, the study provides a comprehensive understanding of how the public perceives the CE and how these perceptions change over time. By drawing data from diverse sources and engaging with these varied segments of society, the study captures a comprehensive view on shift in public sentiment and perspectives related to the CE.
Throughout the process of our study, we meticulously designed and executed a series of systematic experiments with the primary goal of unravelling the intricate relationship between the CE and public sentiment. Our aim was to move beyond mere observation and dive into the underlying factors driving changes in sentiment. To achieve this, we employed topic modelling techniques [10]. To be more specific, we utilised three well-established and open-source topic modelling techniques for their reliability and effectiveness. These include models based on Latent Dirichlet Allocation (LDA) technology [11], Correlation Explanation (CorEx) [12], and BERTopic model [13], which leverages Large Language Model technique on topic modelling. Note that these advanced tools enabled us not only to identify topics related to the CE across various platforms but also to track the ever-evolving and dynamic aspects of public opinion towards CE.
@@
Furthermore, this study explores the complexity of our three models. We have not only implemented training of different models across various datasets but have also fine-tuned the hyperparameters of each model (hyperparameter optimisation). The hyperparameter optimisation process employed a grid search strategy, and optimal hyperparameters were identified for different datasets based on Normalised Pointwise Mutual Information (NPMI) scores [14, 15]. This approach ensures the efficacy and robustness of our analysis. Through the hyperparameter optimisation of the three models, this research found that BERTopic performed best on both The Guardian and Reddit datasets. CorEx achieved the highest NPMI score on Twitter dataset. In Sections 2 and 3, we will discuss the related work and the proposed framework for this study, as well as the experimental settings, respectively. In Section 4, we will comprehensively discuss and analyse each optimal model’s generated topics related to the CE, as well as the evolution of public sentiment towards CE topics over the past few years.
In essence, our research work stands as a pioneering initiative in the exploration of public sentiment across all facets related to the CE. More importantly, we aim to provide valuable insights into the pathways and perspectives guiding us towards a more sustainable future for the CE. We recognise that public sentiment serves as a compass, guiding the trajectory of policies, corporate strategies, and societal attitudes. By shedding light on this intricate relationship, we aim to contribute to the global effort to build a more sustainable CE that benefits us all.
@@ 2 Related work
In this section, we thoroughly examine the extant research and approaches relevant to our study’s scope and objectives. Our study is primarily oriented towards discerning public concerns regarding pivotal aspects of the CE by employing advanced techniques such as topic modelling and the utilisation of large language models. To facilitate a structured and comprehensive review, we categorise the related work into six subsections, each dedicated to an essential facet of our research: Latent Dirichlet Allocation (LDA), Correlation Explanation (CorEx) Topic Model, BERTopic, Evaluation Metrics, and Circular Economy. This section is intended to provide readers with a profound understanding of the foundational research that underpins our investigation.
@@ 2.1 Latent Dirichlet Allocation (LDA)
Understanding the core concepts of LDA is essential for gaining insight into topic modelling. LDA is built upon two fundamental equations, denoted as Equation 1 and Equation 2. For the sake of clarity, we introduce a new variable, denoted as $Q$, which signifies the relationship between each word in each document and the associated topic. $P(W|\alpha,\beta)$ represents the likelihood probability [11] of the entire document collection. The parameter $\alpha$ plays a pivotal role in controlling the mixture of topics within each document, while $\beta$ governs the distribution of words within each topic, determining which words are associated with specific topics. $z_{m,n}$ denotes the topic of the n-th word in the m-th document, $w_{m,n}$ signifies the n-th word in the m-th document, and $\beta_{z_{m,n}}$ represents the vocabulary distribution of topic $z_{m,n}$. In this study, we will not explain the mathematical derivation of LDA in depth. With core Equation 1 and Equation 2 of LDA, we aim is to provide readers with an intuitive understanding of LDA’s operational principles and its robust foundation as a powerful probabilistic model [11].
@@
Among the models designed to address the temporal dimension, the Topics Over Time (TOT) model [18] stands out for its distinctive feature of employing beta-distributed topics, deviating from the convention of Gaussian-distributed topics typically found in Dynamic Topic Models (DTM) [19]. It is noteworthy that the applicability of the TOT model is limited by its constraint of permitting only a point, allowing for one instance of rising and falling trends. This constraint imposes a significant limitation on the model’s ability to capture intricate and multifaceted changes in topics, such as scenarios involving a sequence of topic frequency decline, upswing, and subsequent decline [20]. Our text data exhibits analogous characteristics. Given that our data primarily revolve around the collection of public sentiments and the meticulous analysis of textual content, we anticipate the emergence of fluctuations in topic distributions over time. Hence, we opt for an initial approach utilising the fundamental LDA model to conduct a probabilistic-based topic modelling of the CE. In the study, we employ the fundamental Bag-of-Words (BoW) model [21] as a foundational element in our LDA model. The BoW model focuses exclusively on the presence or absence of individual words within documents, without regard to the sequential order in which they appear. This approach treats phrases such as ‘The Britain loves the circular economy’ and ‘The circular economy loves the Britain’ as equivalent, as the model disregards the specific word order. By discovering the recurrent topics found in the corpus and assessing how common they are in each document, this allows us to establish a foundational understanding of CE topics within the corpus.
Subsequently, we extend our analysis by employing other models: Correlation Explanation (CorEx) Topic Model and BERTopic, offering distinct methodologies for topic modelling. This multi-pronged strategy facilitates a comprehensive exploration of the dynamics surrounding public sentiment and its evolving facets regarding the concept of CE, enabling a deeper comprehension of sentiment changes within the public sentiment.
@@ 2.2 Correlation Explanation (CorEx) Topic Model
where $r(t)$ is an estimate of the marginal distribution of $T$. By maximising this lower bound, CorEx can learn $p(t|x)$, the distribution of topics $T$. CorEx’s strength lies in its semi-supervised nature, enabling the incorporation of prior knowledge to guide the topic discovery process, thereby yielding more interpretable and meaningful topics. In this research, we designated the words ‘reduce’, ‘reuse’, and ‘recycle’ from the 3Rs policy [26] as anchor words (prior knowledge) for the CorEx model, leveraging this prior knowledge to steer the model towards discovering topics pertinent to the circular economy. By employing these anchor words as known topic words, CorEx endeavours to uncover other highly correlated words and topics, better capturing the semantics of the circular economy.
@@ 2.3 BERTopic
BERTopic utilises the pre-trained language model BERT [27] to identify topics in text data. In the study, we employ BERTopic to gain insights into public sentiment regarding the circular economy. The BERTopic with the class-based c-TF-IDF [13] procedure can leverage contextual understanding of words and phrases, and identify topics that may not be apparent through traditional methods [13]. Alamsyah et al. [28] sheds light on the remarkable capabilities of BERTopic in the analysis of sentiment within textual data. In their comparative analysis, BERTopic stands out as an effective model, achieving a notable coherence score [28] of $0.67$ when applied to user comments. This performance is particularly impressive when considering that the text inputs are often concise, aligning with the short text data paradigm, as previously discussed in the Correlation Explanation (CorEx) Topic Model. However, they [13, 28] do not consider the performance of BERTopic on non-short texts (such as our The Guardian news data), nor does it examine whether BERTopic can maintain the same effect on different types of data. As we embark on the study, it is essential to acknowledge that the dataset will encompass text of varying lengths, including non-short text [29]. This diversity in textual data prompts a natural curiosity regarding the adaptability and performance of BERTopic in handling these more extended textual inputs.
@@
UMAP uses the representation of joint probability to represent high-latitude data. McInnes et al. [33] designed a specific family of curves $\frac{1}{1+a(y^{2b})}$ to model the probability of distance in low dimensions, and Equation 7 directly represents the relationship between points $m$ and $n$ in the low dimensions $q_{mn}$. Both $a$ and $b$ in Equation 7 are hyper-parameters in UMAP. In this step, UMAP solves for the hyper-parameters $a$ and $b$ using $min_{d}ist$ a representation of the minimum distance between points in low dimensions. Thus, we can obtain the initial low-dimensional matrix.
@@
In the last step of the whole UMAP algorithm, UMAP will use the Loss function [34] to make the representation of high dimensional points close to the representation of low dimensional points to achieve the effect of dimensional reduction of high-dimensional data. And according to McInnes et al. [33], UMAP will use cross-entropy [35] as the Loss function instead of Kullback Leibler Divergence [36]. This design can reduce the dimensional of the data by preserving the global distance when the data is moved from the high dimensional to the low dimensional [33]. In this paper, we will keep the default setting of UMAP, and cross-Entropy will be used as the loss function of UMAP to realise the dimensionality reduction step of BERTopic. After reducing the data’s dimensional using UMAP, we can employ an appropriate approach (Hierarchical Density-Based Cluster Selection (HDBSCAN) [37] or K-Means [38, 39]) to cluster the data effectively. HDBSCAN is a density clustering algorithm capable of identifying clusters with different shapes and sizes without the need to pre-specifying the number of clusters. K-Means is a clustering algorithm that identifies $K$ clusters within a given dataset. It discovers $K$ distinct clusters, with the centre of each cluster calculated as the mean of the values contained within that cluster. The number of clusters, $K$, is specified by the user, and each cluster is described by its centroid, which is the central point of all the points in the cluster.
@@
From the early experiments, we found that both HDBSCAN and K-Means aided BERTopic in better comprehending the clustering structure of the dimensionality-reduced data. These two methods proved valuable when dealing with more complex and noisy datasets. However, since ensuring the quality of each acquired data point in real-world scenarios is typically challenging and this paper will involve a comparative analysis of three models. Controlling the number of topics generated or clusters formed by different models can provide a clearer understanding of their performance on the same dataset. K-Means allows us to select the desired number of clusters and forces every point to belong to a cluster. Consequently, no outliers will be produced. Nevertheless, as mentioned in BERTopic [13], there are drawbacks. When we forcibly cluster every point, it implies that the cluster is likely to contain noise that may degrade the topic representation. Fortunately, we can leverage the built-in English stop word library in the model to remove unnecessary vocabulary, thereby enhancing the topic representation. All in all, K-Means clustering algorithm will be the choice for this research.
@@
BERTopic modifies the Term Frequency-Inverse document frequency (TF-IDF) algorithm [40]. As the last step of the model, it uses class-based TF-IDF (c-TF-IDF) [13] to extract the topic words. TF-IDF is a statistical technique, same as TF $\cdot$ IDF, wherein TF stands for term frequency, and IDF stands for inverse document frequency. In Equation 8, $i$ represents a specific term, $j$ corresponds to a particular document, and $N$ denotes the total number of words within the document. Equation 8 allows us to compute the term frequency (TF) of a term $i$, which can also be interpreted as the probability of the term $i$ occurring in the content $j$. The IDF for a term $i$ is computed by dividing $N$ by $N_{i}$ which represents the number of documents containing the term $i$ and taking the algorithm of the resulting quotient (Equation 9). However, in cases where term $i$ does not appear in any document, it could lead to a division by zero issue. To prevent this, it is a common practice to $+1$ to the numerator to ensure the formula remains solvable.
From Robertson’s research [40], we understand that TF-IDF is employed to assess the significance of words. When the term frequency (TF) of word $i$ is high in one text but low in other articles, it is deemed that this word $i$ possesses strong category-distinguishing characteristics and is suitable for classification. However, the basic structure of IDF (Inverse Document Frequency) does not take into account the semantic nuances of words or effectively capture word distribution, rendering the IDF’s methodology somewhat rudimentary [41]. Consequently, it struggles to address issues related to word polysemy and homonymy, and as a result, the accuracy of the TF-IDF algorithm is not particularly high, especially when working with pre-classified text datasets [41]. Fortunately, researchers have made advancements in text representations for classification. BERTopic leverages an adapted TF-IDF model, enabling the extraction of keywords based on classified text, which is facilitated through the c-TF-IDF approach [13, 41]. In Equation 10, within the context of the c-TF-IDF algorithm. $W(t,c)$ denotes the weighted score (importance degree) [13] of word $t$ in the category $c$. $f_{t}$ represents the frequency of word $t$ across all categories, and $A$ signifies the average word count in each category. A notable departure from the traditional TF-IDF method can be observed in the IDF formula (Equation 9).

### v2

# Exploring Public Attention in the Circular Economy through Topic Modelling with Twin Hyperparameter Optimisation

@@ Abstract
To advance the circular economy (CE), it is crucial to gain insights into the evolution of public attention, cognitive pathways of the masses concerning circular products, and to identify primary concerns. To achieve this, we collected data from diverse platforms, including Twitter, Reddit, and The Guardian, and utilised three topic models to analyse the data. Given the performance of topic modelling may vary depending on hyperparameter settings, this research proposed a novel framework that integrates twin (single and multi-objective) hyperparameter optimisation for the CE. We conducted systematic experiments to ensure that topic models are set with appropriate hyperparameters under different constraints, providing valuable insights into the correlations between CE and public attention. In summary, our optimised model reveals that public remains concerned about the economic impacts of sustainability and circular practices, particularly regarding recyclable materials and environmentally sustainable technologies. The analysis shows that the CE has attracted significant attention on The Guardian, especially in topics related to sustainable development and environmental protection technologies, while discussions are comparatively less active on Twitter. These insights highlight the need for policymakers to implement targeted education programs, create incentives for businesses to adopt CE principles, and enforce more stringent waste management policies alongside improved recycling processes.
@@ 1 Introduction
The Circular Economy (CE) has gained increased attention from both academics and practitioners in recent years. The primary objectives of the CE are commonly regarded as achieving the twin goals of economic prosperity and improved resource efficiency [1]. However, its impacts on social equity and public perceptions are often overlooked. CE also plays a critical role in transitioning towards net-zero emissions [2] by reducing resource consumption and optimising waste management [3], which lowers energy demand and greenhouse gas emissions [4]. This relationship is especially significant in fields such as energy and artificial intelligence, where integrating advanced technologies and circular principles can lead to substantial improvements in resource efficiency and sustainability.
@@
Moreover, public attention and understanding of CE remain limited. CE represents a paradigm shift from the traditional linear economy—characterised by the ‘take–make–waste’ model [5] to one that promotes the principles of regeneration, material reuse, waste reduction, and pollution prevention. In a world grappling with pressing environmental challenges, public discourse on CE plays a pivotal role in shaping policies [6], business strategies [7], and societal perceptions of sustainable development [8]. Therefore, it has become increasingly crucial to comprehend and interpret public attention concerning the CE. This urgency stems not only from the necessity to understand the intricacies of evolving public attitudes but also from the imperative to analyse the cognitive pathways that influence the public’s perception of recycled products within the broader framework of the CE [9].
Throughout the process of our research, we meticulously designed and executed a series of systematic experiments with the primary goal of unravelling the intricate relationship between CE and public attention. We aimed to move beyond mere observation and dive into the underlying factors driving changes in attention. To achieve this, we employed topic modelling techniques [10]. Topic modelling techniques are a class of machine learning algorithms used to automatically discover the underlying themes or ‘topics’ present in large collections of unstructured text data.
To be more specific, we utilised three well-established and open-source topic modelling techniques for their reliability and effectiveness. These include models based on Latent Dirichlet Allocation (LDA) technology [11], Correlation Explanation (CorEx) [12], and BERTopic model [13]. Using LDA, CorEx, and BERTopic together can provide a more comprehensive understanding of the underlying topics in a dataset, as each technique offers distinct strengths. LDA’s probabilistic model helps identify clear, interpretable topics based on word co-occurrence patterns. CorEx, with its focus on maximising correlations, can discover more nuanced relationships that might not be captured by word frequency alone, leading to more coherent and interpretable topics, especially in complex or highly correlated datasets. BERTopic, which leverages BERT embedding, adds the ability to capture contextual and semantic nuances between words, which is often missed by traditional models like LDA. In this research, we selected three datasets: Twitter, The Guardian, and Reddit for model training. As an old news media, The Guardian has rich official information. In particular, it focuses on the attention of organisations such as governments on CE. Social media platforms such as Twitter and Reddit are rich sources of real-time user-generated content, where people can express their views and ideas on CE. By analysing the data from these platforms, we can capture diverse and contemporary public attention, and thus understand the nuances of public attention to CE around the world.
Furthermore, this research delves into the complexity of the three selected models. We not only trained different models across various datasets but also fine-tuned the hyperparameters of each model using our novel Twin method (both single-objective and multi-objective optimisation techniques). Hyperparameter optimisation is crucial for topic modelling because the performance and quality of the topics generated by models like LDA, CorEx, and BERTopic are sensitive to the selection of hyperparameters. Hyperparameters such as the number of topics can greatly influence the interpretability, coherence, and relevance of the topics. By tuning these hyperparameters, we can ensure that the model effectively captures the underlying structure of the data and produces meaningful, interpretable topics. Without proper optimisation, the model risks either underfitting or overfitting, resulting in irrelevant or overly broad topics. Hyperparameter optimisation explores the hyperparameter space to find optimal configurations, ensuring the best possible performance and more reliable results.
@@
The hyperparameter optimisation processes employed grid search and the Tree-structured Parzen Estimator (TPE), respectively. Optimal hyperparameters were identified for different datasets based on Coherence using Normalised Pointwise Mutual Information ($C_{\text{NPMI}}$) scores [14, 15], as well as diversity [16] and perplexity [10]. This approach ensures the efficacy and robustness of our analysis from multiple criteria. Through the hyperparameter optimisation of the three models, this research found BERTopic performed best on both The Guardian and Reddit datasets. CorEx achieved the highest $C_{\text{NPMI}}$ score on the Twitter dataset.
This research offers several key contributions. First, we collected data from a wide range of sources, including posts from widely used social media platforms (Twitter and Reddit), capturing a wide range of public views on the CE. By analysing the content from The Guardian further reveals distinct differences in how platforms disseminate information and engage in CE discussions. Furthermore, our findings provide a comprehensive view of evolving public perceptions of CE, capturing shifts in attention and attitudes across diverse societal segments. In essence, this research pioneers the exploration of public attention across different facets of the CE, offering valuable insights into the pathways that shape policies, corporate strategies, and societal attitudes towards a sustainable future. This work contributes to the global effort to foster a more sustainable CE.
The rest of the paper is organised as follows. In sections 2 and 3, we will discuss the related work, the proposed framework for this research, and twin hyperparameter optimisation method, as well as the experimental settings, respectively. In Section 4, we will comprehensively discuss and analyse each optimal model’s generated topics related to the CE, as well as the evolution of public attention towards CE topics over the past few years. Section 5 offers concluding remarks, policy implications and pinpoints potential avenues for future studies.
@@ 2 Related work
In this section, we thoroughly examine the extant research and approaches relevant to our research’s scope and objectives. Our research is primarily oriented towards discerning public concerns regarding pivotal aspects of the CE by employing advanced techniques such as topic modelling and the utilisation of language models. To facilitate a structured and comprehensive review, we categorise the related work into six subsections, each dedicated to an essential facet of our research: Latent Dirichlet Allocation (LDA), Correlation Explanation (CorEx) Topic Model, BERTopic, Evaluation Metrics, and Circular Economy. This section is intended to provide readers with a profound understanding of the foundational research that underpins our investigation.
@@ 2.1 Latent Dirichlet Allocation (LDA)
Understanding the core concepts of LDA is essential for gaining insight into topic modelling. LDA is built upon two fundamental equations, denoted as Equation 1 and Equation 2. For the sake of clarity, we introduce a new variable, denoted as $Q$, which signifies the relationship between each word in each document and the associated topic. $P(W|\alpha,\beta)$ represents the likelihood probability [11] of the entire document collection. The parameter $\alpha$ plays a pivotal role in controlling the mixture of topics within each document, while $\beta$ governs the distribution of words within each topic, determining which words are associated with specific topics. $z_{m,n}$ denotes the topic of the n-th word in the m-th document, $w_{m,n}$ signifies the n-th word in the m-th document, and $\beta_{z_{m,n}}$ represents the vocabulary distribution of topic $z_{m,n}$. In this research, we will not explain the mathematical derivation of LDA in depth. With core Equation 1 and Equation 2 of LDA, we aim to provide readers with an intuitive understanding of LDA’s operational principles and its robust foundation as a powerful probabilistic model [11].
@@
Among the models designed to address the temporal dimension, the Topics Over Time (TOT) model [19] stands out for its distinctive feature of employing beta-distributed topics, deviating from the convention of Gaussian-distributed topics typically found in Dynamic Topic Models (DTM) [20]. It is noteworthy that the applicability of the TOT model is limited by its constraint of permitting only a point, allowing for one instance of rising and falling trends. This constraint imposes a significant limitation on the model’s ability to capture intricate and multifaceted changes in topics, such as scenarios involving a sequence of topic frequency decline, upswing, and subsequent decline [21]. Our text data exhibits analogous characteristics. Given that our data primarily revolve around the collection of public attention and the meticulous analysis of textual content, we anticipate the emergence of fluctuations in topic distributions over time. Hence, we opt for an initial approach utilising the fundamental LDA model to conduct a probabilistic-based topic modelling of the CE. In the research, we employ the fundamental Bag-of-Words (BoW) model [22] as a foundational element in our LDA model. The BoW model focuses exclusively on the presence or absence of individual words within documents, without regard to the sequential order in which they appear. This approach treats phrases such as ‘Britain loves the circular economy’ and ‘The circular economy loves Britain’ as equivalent, as the model disregards the specific word order. By discovering the recurrent topics found in the corpus and assessing how common they are in each document, allows us to establish a foundational understanding of CE topics within the corpus.
Subsequently, we extend our analysis by employing other models: Correlation Explanation (CorEx) Topic Model and BERTopic, offering distinct methodologies for topic modelling. This multi-pronged strategy facilitates a comprehensive exploration of the dynamics surrounding public attention and its evolving facets regarding the concept of CE, enabling a deeper comprehension of attention changes within public attention.
@@ 2.2 Correlation Explanation (CorEx) Topic Model
where $r(t)$ is an estimate of the marginal distribution of $T$. By maximising this lower bound, CorEx can learn $p(t|x)$, the distribution of topics $T$. CorEx’s strength lies in its semi-supervised nature, enabling the incorporation of prior knowledge to guide the topic discovery process, thereby yielding more interpretable and meaningful topics. In this research, we designated the words ‘reduce’, ‘reuse’, and ‘recycle’ from the 3Rs policy [27] as anchor words (prior knowledge) for the CorEx model, leveraging this prior knowledge to steer the model towards discovering topics pertinent to the circular economy. By employing these anchor words as known topic words, CorEx endeavours to uncover other highly correlated topics and better capture the semantics of the circular economy.
@@ 2.3 BERTopic
BERTopic utilises the pre-trained language model BERT [28] to identify topics in text data. In the research, we employ BERTopic to gain insights into public attention regarding the circular economy. The BERTopic with the class-based c-TF-IDF [13] procedure can leverage contextual understanding of words and phrases, and identify topics that may not be apparent through traditional methods [13]. Alamsyah et al. [29] sheds light on the remarkable capabilities of BERTopic in the analysis of attention within textual data. In their comparative analysis, BERTopic stands out as an effective model, achieving a notable coherence score [29] of $0.67$ when applied to user comments. This performance is particularly impressive when considering that the text inputs are often concise, aligning with the short text data paradigm, as previously discussed in the Correlation Explanation (CorEx) Topic Model. However, they [13, 29] do not consider the performance of BERTopic on non-short texts (such as our The Guardian news data), nor does it examine whether BERTopic can maintain the same effect on different types of data. As we embark on the research, it is essential to acknowledge that the dataset will encompass text of varying lengths, including non-short text [30]. This diversity in textual data prompts a natural curiosity regarding the adaptability and performance of BERTopic in handling these more extended textual inputs.
@@
UMAP uses the representation of joint probability to represent high-latitude data. McInnes et al. [34] designed a specific family of curves $\frac{1}{1+a(y^{2b})}$ to model the probability of distance in low dimensions, and Equation 7 directly represents the relationship between points $m$ and $n$ in the low dimensions $q_{mn}$. Both $a$ and $b$ in Equation 7 are hyperparameters in UMAP. In this step, UMAP solves for the hyperparameters $a$ and $b$ using $min_{d}ist$ a representation of the minimum distance between points in low dimensions. Thus, we can obtain the initial low-dimensional matrix.
@@
In the last step of the whole UMAP algorithm, UMAP will use the Loss function [35] to make the representation of high dimensional points close to the representation of low dimensional points to achieve the effect of dimensional reduction of high-dimensional data. And according to McInnes et al. [34], UMAP will use cross-entropy [36] as the Loss function instead of Kullback Leibler Divergence [37]. This design can reduce the dimensional of the data by preserving the global distance when the data is moved from the high dimensional to the low dimensional [34]. In this paper, we will keep the default setting of UMAP, and cross-entropy will be used as the loss function of UMAP to realise the dimensionality reduction step of BERTopic. After reducing the data’s dimensional using UMAP, we can employ an appropriate approach (Hierarchical Density-Based Cluster Selection (HDBSCAN) [38] or K-Means [39, 40]) to cluster the data effectively. HDBSCAN is a density clustering algorithm capable of identifying clusters with different shapes and sizes without the need to pre-specify the number of clusters. K-Means is a clustering algorithm that identifies $K$ clusters within a given dataset. It discovers $K$ distinct clusters, with the centre of each cluster calculated as the mean of the values contained within that cluster. The number of clusters, $K$, is specified by the user, and each cluster is described by its centroid, which is the central point of all the points in the cluster.
@@
From the early experiments, we found that both HDBSCAN and K-Means aided BERTopic in better comprehending the clustering structure of the dimensionality-reduced data. These two methods proved valuable when dealing with more complex and noisy datasets. However, since ensuring the quality of each acquired data point in real-world scenarios is typically challenging and this paper will involve a comparative analysis of three models. Controlling the number of topics generated or clusters formed by different models can provide a clearer understanding of their performance on the same dataset. K-Means allows us to select the desired number of clusters and forces every point to belong to a cluster. Consequently, no outliers will be produced. Nevertheless, as mentioned in BERTopic [13], there are drawbacks. When we forcibly cluster every point, it implies that the cluster is likely to contain noise that may degrade the topic representation. Fortunately, we can leverage the built-in English stop word library in the model to remove unnecessary vocabulary, thereby enhancing the topic representation. All in all, the K-Means clustering algorithm will be the choice for this research.
@@
BERTopic modifies the Term Frequency-Inverse document frequency (TF-IDF) algorithm [41]. As the last step of the model, it uses class-based TF-IDF (c-TF-IDF) [13] to extract the topic words. TF-IDF is a statistical technique, the same as TF $\cdot$ IDF, wherein TF stands for term frequency, and IDF stands for inverse document frequency. In Equation 8, $i$ represents a specific term, $j$ corresponds to a particular document, and $N$ denotes the total number of words within the document. Equation 8 allows us to compute the term frequency (TF) of a term $i$, which can also be interpreted as the probability of the term $i$ occurring in the content $j$. The IDF for a term $i$ is computed by dividing $N$ by $N_{i}$ which represents the number of documents containing the term $i$ and taking the algorithm of the resulting quotient (Equation 9). However, in cases where the term $i$ does not appear in any document, it could lead to a division by zero issue. To prevent this, it is a common practice to $+1$ to the numerator to ensure the formula remains solvable.
From Robertson’s research [41], we understand that TF-IDF is employed to assess the significance of words. When the term frequency (TF) of the word $i$ is high in one text but low in other articles, it is deemed that this word $i$ possesses strong category-distinguishing characteristics and is suitable for classification. However, the basic structure of IDF (Inverse Document Frequency) does not take into account the semantic nuances of words or effectively capture word distribution, rendering the IDF’s methodology somewhat rudimentary [42]. Consequently, it struggles to address issues related to word polysemy and homonymy, and as a result, the accuracy of the TF-IDF algorithm is not particularly high, especially when working with pre-classified text datasets [42]. Fortunately, researchers have made advancements in text representations for classification. BERTopic leverages an adapted TF-IDF model, enabling the extraction of keywords based on classified text, which is facilitated through the c-TF-IDF approach [13, 42]. In Equation 10, within the context of the c-TF-IDF algorithm. $W(t,c)$ denotes the weighted score (importance degree) [13] of word $t$ in the category $c$. $f_{t}$ represents the frequency of word $t$ across all categories, and $A$ signifies the average word count in each category. A notable departure from the traditional TF-IDF method can be observed in the IDF formula (Equation 9).